OpenStudy (anonymous):

Tori's 8-ounce glass of soymilk has 6 grams of protein. Brady's 16-ounce glass has 12 grams of protein. What is ratio of Tori's drink size to Brady's drink size?

OpenStudy (anonymous):

plz help mee quickly!!

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

thats not one of my answers

OpenStudy (anonymous):

choices*

OpenStudy (anonymous):

@Satellite73 its either going to be : 1:1, 4:3, 16:4, 3:4.

OpenStudy (anonymous):

then there is a mistake in your answers because \[\frac{8}{16}=\frac{1}{2}\]

OpenStudy (anonymous):

ratio of size of drink to protein amount is \[\frac{8}{6}=\frac{4}{3}\] but that was not the question asked